# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

### v1

@@ 3 Results
Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions. In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 8). However extrapolation shows this model is not theoretically correct (Figure 13 in SI).
This dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92]. Because this expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search. We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop). We found that when using basic operators and context, GPT-4 was getting close to finding the target model (Figure 8) upon running 15 iterations with this feedback loop. To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed. Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR. But when targeting models like dual-site Langmuir, these instructions may have been a liability, and we needed to adjust the prompt and feedback sent to GPT-4 to allow for the exploration of more complex expressions.
@@
Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to GPT-4 exceeded the token limit. Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations. So, we developed a cost-saving scheme: only send a portion of the data in the prompt to GPT-4, while fitting and evaluating the generated expressions using the whole dataset (Figure 14 in SI).
@@
Since GPT-4 generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging. We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions. So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error. Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to GPT-4 for feedback. To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
@@
Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6]. We modified the basic prompt to encourage GPT-4 to explore longer expressions and also tested the effect of “prodding" GPT-4 by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better. We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) random data points and the second (S2) with another random 36 (20%) data points. Overall, feeding in more data generated better-fitted and more complex expressions. Figure 9 shows the top expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^).
@@
Figure 9: Models for Nikuradse dataset from GPT-4. Here P refers to prompt versions, and S refers to dataset. We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig. 9 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways. However it generalizes better with fewer data for prompt version 3 where we challenge GPT-4 by providing information about the MAE and complexity of BMS model.
@@
We found an optimal expression from GPT-4 with a complexity of 41 and an MAE of 0.01086 (Figure 9 C). The MAE is approximately three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS). BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392. However unlike GPT-4, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps to identify this expression. Our GPT-4 model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions. We additionally assessed BMS on its default move probabilities and using 40 temperatures for parallel tempering as mentioned in the paper, and ran it for 153 MC steps with Nikuradse data -- it generated a constant function, and upon running it longer (1000 MC steps), it suggests an expression with MAE 0.13436 and complexity 25. Thus, we can see that GPT-4’s incremental suggestions for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS. However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost. Using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
@@ 4 Discussion and Conclusion
SR programs that optimize for speed aim to generate expressions quickly. In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively. The clearest cases of “leveraging the context” occurred when GPT-4’s first guess included the target model among three expressions. But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions. However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context. Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens.
@@
In general, we found natural language to be a rather clumsy interface for controlling expression length. Different prompts and feedback mechanisms led to distributions of expressions with varying length. Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [93]. We ran separate, focused tests to evaluate the effect of prompting on expression length, and GPT-4 did not obey instructions that requested, for example, "expressions of length 17." Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length. The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting complexity.
@@
Testing our approach on equation rediscovery using GPT-4 invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources. Indeed, the data are publicly available, as well, though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets. We found strong evidence of this because of our scratchpad implementation, which revealed when it was thinking of Kepler’s law before seeing the data. Nonetheless, we found evaluating the model outputs to be informative.
@@
Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include. This context may include experimental details, instrument specifications, and literature, and even a “memorized” explanation of the data by an LLM or retrieved from context in a retrieval-augmented generation scheme [94] would be valuable. A true blind test would be to rediscover a novel scientific law using an LLM with a knowledge cutoff predating the seminal publication.
@@ Acknowledgements
The code, prompts, and data supporting the findings of this work are available at https://github.com/ATOMSLab/pySR_adsorption.
@@ 5 Supporting Information
Figure 11: Context provided to GPT-4 for all the datasets
Figure 12: Datasets explored using GPT-4 for SR

### v2

@@ 3 Results
Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions. In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 12). However, extrapolation shows this model is not theoretically correct (Figure 19 in SI). This may also explain why we see a decreasing score trend, even after identifying the target expression, compared to other datasets (Figure 13). We analyzed the LLM-generated expressions from both dual-site and Langmuir adsorption datasets by checking whether they satisfy thermodynamic constraints by running through a computer algebra system built in our previous work [28]. We consistently found one more accurate expression at the expense of an added complexity for Henry’s Law term, which models adsorption at low pressures well.
@@
Dual-site Langmuir dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92]. Because the target model expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search in GPT-4 model. We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop). We found that when using basic operators (easy search) and all tools on, GPT-4 was getting close to finding the target model (Figure 12) upon running 15 iterations with this feedback loop. To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed. Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR.
Figure 13: GPT-4 and GPT-4o on dual-site Langmuir’s adsorption data. All the trends in GPT-4 model overlap at a score of 0. Transparency and different marker styles are used to differentiate these trends.
However, the same prompts were more successful with GPT-4o model, which reasoned better with data and context. For experiments without data, GPT-4o relied on the context of general adsorption behavior. It initiated with expressions based on Langmuir, Freundlich, and occasionally Temkin and BET theories. Interestingly, we observed the highest score on this dataset when no context was provided, and the target model was rediscovered through dataset patterns and iteration feedback. Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in the datasets related to saturation behavior and leading to modifications of the Langmuir expression to derive the dual-site behavior.
Figure 14: Instances of GPT-4o reasoning on dual-site Langmuir dataset. The top snippet of the scratchpad is from Run 1, and the bottom is from Run 2. Both are for runs with all tools on and iteration 4.
@@
Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to LLMs exceeded the token limit. Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations. So, we developed a cost-saving scheme: only send a portion of the data in the prompt to LLM, while fitting and evaluating the generated expressions using the whole dataset (Figure 20 in SI).
@@
Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging. We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions. So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error. Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to LLMs for feedback. To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
@@
Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6]. We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better. We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) data points and the second (S2) with another 36 (20%) data points. Overall, feeding in more data generated better-fitted and more complex expressions. Figure 15 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^). We also list the complexities and MAEs for these expressions in Table 1 in SI.
@@
Figure 15: Models for Nikuradse dataset. Here, P refers to prompt versions and S refers to dataset. The expressions generated by the GPT-4o model are marked with an ‘o’ at the end. We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig. 15 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways. However, GPT-4o model only shows unphysical behavior in prompt version 1 with fewer data. Both the models generalize better with fewer data in prompt version 3 where we challenge by providing information about the MAE and complexity of the BMS model.
@@
We found optimal expressions from GPT-4 and GPT-4o with a complexity of 41 and 27 and an MAE of 0.01086 and 0.00924, respectively (Figure 15 C). GPT-4o performs qualitatively better, generating a more accurate expression at a significantly lower complexity than GPT-4. The MAE from both models is at least two to three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS). BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392. However unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression. Our model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions. The optimal expression chosen is the best from 5 runs (1/153*5 or 1/765). The incremental suggestions from LLMs for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS. However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost. For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
@@ 4 Discussion and Conclusions
SR programs that optimize for speed aim to generate expressions quickly. In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively. The clearest cases of “leveraging the context” occurred when the first guess included the target model among three expressions. But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions. However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context. Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens. GPT-4o is an upgraded version of GPT-4 that offers improved performance, being both faster and more cost-effective [93]. Consistent with observations made by others who have tested it on various tasks [94, 95], we also noted improvements in the quality of generated expressions and the effectiveness of scratchpad analysis through reasoning for symbolic regression.
@@
In general, we found natural language to be a rather clumsy interface for controlling expression length. Different prompts and feedback mechanisms led to distributions of expressions with varying length. Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [96]. We ran separate, focused tests to evaluate the effect of prompting on expression length with the GPT-4 model. However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length. The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting operators and variables to calculate complexity.
@@
Testing our approach on equation rediscovery using language models invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources. Indeed, the data are publicly available, as well (appearing, for example, as tables in the .pdf documents of the historical papers), though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets. We found strong evidence of this when the scratchpad revealed the LLM naming Kepler’s law before seeing the data. Nonetheless, we found evaluating the model outputs to be informative.
@@
Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include. Our experiments included $<$100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [97].
The recent study by Shojaee and coworkers [37] demonstrated strong performance in generating scientific expressions using LLMs, using an approach that generated more structured outputs than our approach. Rather than rediscovering scientific laws as standalone expressions, they tasked the LLMs to “autocomplete” a structured Python function within the context of an optimization problem. They also designed new benchmark scenarios and datasets to highlight the strengths of this approach. This approach effectively leverages the strengths of language models in reasoning about well-defined problem statements. However, the extent to which its success can be attributed to prior exposure to similar questions during training remains unclear without a scratchpad to gain insights into its reasoning steps. A true blind test with these language models would be to rediscover an unpublished or recently-published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
@@ Acknowledgements
The code, prompts, and data supporting the findings of this work are available at https://github.com/ATOMSLab/LLMsforSR.
@@ 5 Supporting Information
Figure 17: Context provided to LLMs for all the datasets
Figure 18: Datasets explored using LLMs for SR
@@
Expression Mean Absolute Error (MAE) Complexity
P1S1 $0.02270419$ $13$
P1S2 $0.00978477$ $29$
P2S1 $0.00897093$ $69$
P2S2 $0.00931620$ $49$
P3S1 $0.01086397$ $41$
P3S2 $0.00992712$ $49$
P1S1o $0.02007803$ $19$
P1S2o $0.02141686$ $17$
P2S1o $0.00954461$ $27$
P2S2o $0.01186963$ $27$
P3S1o $0.00923655$ $27$
P3S2o $0.01144178$ $19$
Table 1: Expressions of LLM models for Nikuradse dataset with complexity and MAE. Here, P refers to prompt versions and S refers to dataset. The expressions generated by the GPT4o model are marked with an ‘o’ at the end.